# WorldForge: Unlocking Emergent 3D/4D Generation in Video Diffusion Model via Training-Free Guidance

### v1

@@ 3.1.1 Denoising Diffusion Models and Guidance
Diffusion Solvers. Modern generative modeling is dominated by two paradigms: diffusion models (Ho et al., 2020; Song et al., 2020a) and flow-based models (Lipman et al., 2022). Under the SDE view, diffusion models admit a deterministic ODE limit that connects them to flow-based formulations through reparameterization (Gao et al., 2025). The detailed derivation of this equivalence is provided in Appendix A. In this section, we take diffusion models as representative examples. Consider the widely used DDIM sampler (Song et al., 2020a): it recovers the clean sample $\mathbf{x}_{0}$ by reversing the forward noising of a Gaussian prior $\mathbf{x}_{T}$. Given a noise-prediction network $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, the sampler estimates an intermediate signal $\hat{\mathbf{x}}_{0}$ from the current state $\mathbf{x}_{t}$:
@@
Iterating this update from $t=T$ to $t=0$ produces the final sample $\mathbf{x}_{0}$. Our method intervenes at this stage by modifying $\hat{\mathbf{x}}_{0}$ to enforce trajectory control. Notably, other popular solvers, such as UniPC, EDM, and PNDM (Zhao et al., 2023; Karras et al., 2022; Liu et al., 2022a), also compute $\hat{\mathbf{x}}_{0}$ directly or can recover it via a parameterized transformation, so our framework is broadly compatible.
Classifier Free Guidance. To improve fidelity to the condition, CFG (Ho & Salimans, 2021) adjusts the score function during sampling:
$\tilde{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},t)=\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\phi)+\omega_{\text{CFG}}\cdot\left[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\phi)\right],$ (3)
where $\omega_{\text{CFG}}$ is the guidance weight, with $\mathbf{c}$ and $\phi$ denoting the conditional and unconditional branches, respectively. This interpolates conditional and unconditional scores to steer the sampling trajectory. Our approach extends this principle through a self-referential guidance mechanism that dynamically adjusts the guided prediction using the model’s own unguided output at each step.
@@ 3.1.2 Trajectory Control via Depth-Based Warping
Our framework uses a warping-and-repainting strategy, embedding geometric cues through depth-based view warping. For reliable depth estimation to support warping into new views, we employ depth prediction networks that takes one or more input images $\{\mathbf{I}_{i}\}_{i=1}^{N}$ and estimates corresponding camera poses and depth maps:
$f:\{\mathbf{I}_{i}\}_{i=1}^{N}\rightarrow\{\mathbf{P}_{i},\mathbf{D}_{i}\},$ (4)
where the depth maps $\mathbf{D}_{i}$ and poses $\mathbf{P}_{i}$ are then used to warp the source views to target poses. Formally, the warping function $\mathcal{W}$ projects source frames $\mathbf{I}_{src}$ with depth $\mathbf{D}_{src}$ from pose $\mathbf{P}_{src}$ to target pose $\mathbf{P}_{tar}$, producing partial target views $\mathbf{I}^{\prime}_{tar}$ and validity masks $\mathbf{M}_{tar}$, which indicate the visible pixels in warped views:
@@
The warped frames $\mathbf{I}^{\prime}_{tar}$ and masks $\mathbf{M}_{tar}$ provide trajectory-aware observations that serve as the basis for generating videos along arbitrary target poses $\mathbf{P}_{tar}$, though limited to regions visible in source views.
With these preliminaries, we now present our method. Our goal is to use the trajectory control introduced in this section to guide video generation in VDM. To achieve precise and consistent control, we design three modules: Intra-Step Recursive Refinement, Flow-Gated Latent Fusion, and Dual-Path Self-Corrective Guidance. Together, they complement each other to ensure accurate trajectory guidance while preserving high-quality synthesis.
@@ 3.2 Intra-Step Recursive Refinement
To enable precise trajectory injection during VDM’s inference processing, we introduce Intra-Step Recursive Refinement (IRR). As noted in Sec. 3.1.1, the denoising process produces an intermediate variable $\hat{\mathbf{x}}_{0}^{(i)}$, a coarse estimate of the final output and the baseline for later steps. Building on this, IRR modifies $\hat{\mathbf{x}}_{0}^{(i)}$ to impose trajectory constraints, ensuring that generation follows the desired path.
IRR operates within the update process of Eq. (1) and Eq. (2). Given the one-step denoised sample $\hat{\mathbf{x}}_{0}^{(i)}$ from Eq. (1), we fuse it with the trajectory latent $\mathbf{Z}_{\text{traj}}$, obtained by encoding the warped frames from Eq. (5) into the latent space. We then add small Gaussian noise $\bm{\epsilon}$ to obtain a modified latent $\mathbf{x}_{t_{i}}^{\prime}$:
$\mathbf{x}_{t_{i}}^{\prime}=(1-w(\sigma))\,\mathbf{F}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})+w(\sigma)\,\cdot\bm{\epsilon},$ (6)
where $\bm{\epsilon}=\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the initial Gaussian noise used in sampling. To return the fused result $\mathbf{x}_{t_{i}}^{\prime}$ into the denoising process while injecting trajectory information, we reintroduce noise $\bm{\epsilon}$. Here $0<w(\cdot)<1$ is a user-defined weighting factor derived from the noise schedule $\sigma$, which controls the strength of the reintroduced noise. $\mathbf{F}(\cdot)$ is a mask-based fusion function defined as $\mathbf{F}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})=\mathbf{M}\cdot\mathbf{Z}_{\text{traj}}+(1-\mathbf{M})\cdot\hat{\mathbf{x}}_{0}^{(i)},$ where $\mathbf{M}$ is the binary validity mask from Eq. (5). It copies observable warped content from $\mathbf{Z}_{\text{traj}}$ into the corresponding locations of $\hat{\mathbf{x}}_{0}^{(i)}$, while leaving unobserved regions unchanged.
In summary, IRR embeds a micro predictor–corrector at each denoising step. By updating $\hat{\mathbf{x}}_{0}^{(i)}$ with explicit trajectory cues, it continually corrects the sampling path and ensures that synthesis follows the target trajectory precisely.
@@ 3.3 Flow-Gated Latent Fusion
In the IRR process, directly overwriting all latent channels with trajectory information often degrades visual quality. This is because VAE latents tend to encode different types of information: some channels mainly capture appearance, while others relate more to motion. Using Eq. (6) to replace all channels indiscriminately introduces noise into appearance-dominant ones. To avoid this, we introduce Flow-Gated Latent Fusion (FLF), which decouples motion and appearance features in the latent space and selectively injects trajectory guidance into motion-relevant channels.
To separate motion from appearance and identify motion-relevant channels, we introduce a flow-based scoring scheme. Optical flow captures pixel-wise motion between frames and is widely used to describe temporal dynamics. By comparing predicted and reference flows, we estimate how strongly each latent channel encodes motion.
Given the predicted latent $\hat{\mathbf{x}}_{0}^{(i)}$ from IRR and the reference trajectory latent $\mathbf{Z}_{\text{traj}}$, we compute per-channel flows using the Farnebäck algorithm. For each channel $c$, we obtain predicted flow $\mathcal{F}^{(c)}_{\text{pred}}(x,y,\tau)=(u^{(c)}_{\text{pred}}(x,y,\tau),v^{(c)}_{\text{pred}}(x,y,\tau))$ and reference ground truth (GT) flow $\mathcal{F}^{(c)}_{\text{gt}}(x,y,\tau)=(u^{(c)}_{\text{gt}}(x,y,\tau),v^{(c)}_{\text{gt}}(x,y,\tau))$, restricted to valid regions $\mathbf{M}^{(c)}(x,y,\tau)$ from Eq. (5). Here $(x,y)$ are pixel coordinates, $\tau$ is the time index, and $u_{*},v_{*}$ are the horizontal and vertical components.
To evaluate motion relevance, we design a normalized score that combines three masked metrics: Masked End-point Error (M-EPE), Masked Angular Error (M-AE), and Outlier Percentage (Fl-all). M-EPE is the average distance between the vectors of the predicted flow and GT flow; smaller values mean better alignment. M-AE measures the angular difference between flows; smaller values mean more accurate directions. Fl-all is the percentage of pixels with large deviations; lower values mean fewer outliers and higher reliability. Together, these metrics provide a robust estimate of each channel’s motion relevance. The formulas are given below:
M-EPE: Given the set $\Omega$ of all valid pixels $(x,y,\tau)$ where $\mathbf{M}^{(c)}=1$, we compute the per-channel $L_{2}$ distance between predicted and GT flows and normalize by the number of valid pixels $|\Omega|$:
$\text{M-EPE}_{c}=\frac{1}{|\Omega|}\sum_{(x,y,\tau)\in\Omega}\left\|\mathcal{F}^{(c)}_{\text{pred}}(x,y,\tau)-\mathcal{F}^{(c)}_{\text{gt}}(x,y,\tau)\right\|_{2}.$ (7)
M-AE: Similar to M-EPE, for each valid pixel in $\Omega$ we compute the cosine similarity between $\mathcal{F}^{(c)}_{\text{pred}}$ and $\mathcal{F}^{(c)}_{\text{gt}}$, and use it to measure angular error:
$\text{M-AE}_{c}=\frac{1}{|\Omega|}\sum_{(x,y,\tau)\in\Omega}\arccos\left(\frac{\mathcal{F}^{(c)}_{pred}(x,y,\tau)\cdot\mathcal{F}^{(c)}_{gt}(x,y,\tau)}{\|\mathcal{F}^{(c)}_{pred}(x,y,\tau)\|\cdot\|\mathcal{F}^{(c)}_{gt}(x,y,\tau)\|}\right).$ (8)
Fl-all: A valid pixel in $\Omega$ is an outlier if its M-EPE exceeds $3$ or its relative error exceeds $5\%$. Fl-all (denoted as $\text{F}_{c}$) is the percentage of outliers among valid pixels; lower values mean higher reliability.
To unify the scales of the three metrics and simplify score computation, we normalize $\text{M-EPE}_{c},\text{M-AE}_{c},\text{F}_{c}$ to $[0,1]$, denoted as $\text{E}^{\text{Norm}}_{c},\text{A}^{\text{Norm}}_{c},$ and $\text{F}^{\text{Norm}}_{c}$, respectively:
$\text{E}^{\text{Norm}}_{c}=\min(\text{M-EPE}_{c}/n_{1},1),\quad\text{A}^{\text{Norm}}_{c}=\min(\text{M-AE}_{c}/n_{2},1),\quad\text{F}^{\text{Norm}}_{c}=\min(\text{F}_{c}/n_{3},1),$ (9)
where $n_{1},n_{2},n_{3}$ are normalization factors. We then define the final flow score for each channel as $S$, which directly reflects the consistency between the predicted and GT flows:
$S^{(c)}=\gamma_{1}\!\left(1-\text{E}^{\text{Norm}}_{c}\right)+\gamma_{2}\!\left(1-\text{A}^{\text{Norm}}_{c}\right)+\gamma_{3}\!\left(1-\text{F}^{\text{Norm}}_{c}\right),$ (10)
where $\gamma_{1}+\gamma_{2}+\gamma_{3}=1$. The flow score $S^{(c)}$ measures the motion relevance of channel $c$. A higher score means its predicted flow is closer to the ground truth and thus encodes more motion. A lower score indicates misalignment and suggests that the channel encodes more non-motion information, such as appearance. This score enables us to decouple motion and appearance channels in the latent space. To select motion-relevant channels, we set a dynamic threshold:
$\delta^{(i)}=\mu_{S}^{(i)}-\lambda^{(i)}\sigma_{S}^{(i)},$ (11)
where $\mu_{S}^{(i)}$ and $\sigma_{S}^{(i)}$ are the mean and standard deviation of $\{S^{(i,c)}\}$ at step $i$. Channels with scores above $\delta^{(i)}$ are regarded as motion-relevant and only these channels receive trajectory guidance. The selection sets are:
$\mathcal{C}_{\text{sel}}^{(i)}=\{\,c\mid S^{(i,c)}\geq\delta^{(i)}\,\}.$
Only these channels are updated with trajectory latents $\mathbf{Z}_{\text{traj}}$. The parameter $0<\lambda^{(i)}<1$ controls the strictness of selection and decreases during denoising. Early steps therefore replace more channels for stronger trajectory adherence, while later steps replace fewer channels to preserve appearance details. Finally, the latent update rule for $\hat{\mathbf{x}}_{0}^{(i)}$ is:
@@
$\mathbf{FLF}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})^{(c)}=\begin{cases}\mathbf{M}^{(c)}\cdot\mathbf{Z}_{\text{traj}}^{(c)}+\left(1-\mathbf{M}^{(c)}\right)\cdot\hat{\mathbf{x}}_{0}^{(i,c)},&c\in\mathcal{C}_{\text{sel}}^{(i)},\\[2.0pt] \hat{\mathbf{x}}_{0}^{(i,c)},&\text{otherwise}.\end{cases}$ (12)
where $\mathbf{FLF}(\cdot)$ denotes the flow-score–based selective fusion. We use it to replace $\mathbf{F}(\cdot)$ in Eq. (6), which overwrote all channels without distinction. The updated fusion rule is:
$\mathbf{x}_{t_{i}}^{\prime}=(1-w(\sigma))\,\mathbf{FLF}(\hat{\mathbf{x}}_{0}^{(i)},\mathbf{Z}_{\text{traj}})+w(\sigma)\,\cdot\bm{\epsilon},$ (13)
In summary, FLF enables fine-grained trajectory injection while preserving model priors and synthesis quality. Unlike Restart Sampling (Xu et al., 2023), which restarts the schedule, or ViewExtrapolator (Liu et al., 2024), which injects trajectory data directly into $\hat{\mathbf{x}}_{0}^{(i)}$, our method uses IRR to apply fine-grained guidance at every timestep. Combined with FLF, it separates motion from appearance and delivers precise, high-quality trajectory control at inference.
@@ 3.4 Dual-Path Self-Corrective Guidance
Trajectory latents $\mathbf{Z}_{\text{traj}}$ obtained by warping along the target motion often suffer from distortions due to depth errors, occlusions, or misalignments. Such artifacts are common in warp-based methods and degrade synthesis quality. To address this, we draw inspiration from CFG (Ho & Salimans, 2021), which interpolates between conditional and unconditional predictions. Instead of using conditioning, we exploit the discrepancy between two denoising paths in IRR to obtain a more reliable direction. Based on this idea, we propose Dual-Path Self-Corrective Guidance (DSG). At each iteration, IRR produces two velocity fields: the unguided $\mathbf{v}_{t_{i}}^{\text{ori}}$ from the original latent $\mathbf{x}_{t_{i}}$, and the guided $\mathbf{v}_{t_{i}}^{\text{traj}}$ from the corrected latent $\mathbf{x}^{\prime}_{t_{i}}$ obtained by injecting trajectory cues. DSG leverages their difference to form a correction signal that steers denoising toward a higher-quality path.
At each step $t_{i}$, IRR and FLF generate two velocity fields: $\mathbf{v}_{t_{i}}^{\text{ori}}$ and $\mathbf{v}_{t_{i}}^{\text{traj}}$. The guided velocity $\mathbf{v}_{t_{i}}^{\text{traj}}$ enforces trajectory adherence but may add noise from imperfect warping, while the unguided velocity $\mathbf{v}_{t_{i}}^{\text{ori}}$ stays on the data manifold with higher fidelity but ignores trajectory cues. To exploit these two directions for a better denoising path, we propose DSG, which combines them to balance control and fidelity. Unlike CFG (Eq. (3)) (Ho & Salimans, 2021), which interpolates unconditional and conditional predictions, DSG uses two trajectory-specific complementary branches designed for motion control. The corrected velocity $\mathbf{v}_{t_{i}}^{\text{corr}}$ is then computed as:
@@
$\mathbf{v}_{t_{i}}^{\text{corr}}=\mathbf{v}_{t_{i}}^{\text{traj}}+\rho\cdot\beta_{t_{i}}\bigl(\mathbf{v}_{t_{i}}^{\text{traj}}-\mu\cdot\alpha_{t_{i}}\cdot\mathbf{v}_{t_{i}}^{\text{ori}}\bigr),$ (14)
where $\rho$ controls guidance strength, $\mu$ normalizes magnitudes, $\alpha_{t_{i}}=(\mathbf{v}_{t_{i}}^{\text{traj}}\cdot\mathbf{v}_{t_{i}}^{\text{ori}})/(\lVert\mathbf{v}_{t_{i}}^{\text{traj}}\rVert\cdot\lVert\mathbf{v}_{t_{i}}^{\text{ori}}\rVert)$ is the cosine similarity between guided and unguided velocities, and $\beta_{t_{i}}=\sin(\arccos(\alpha_{t_{i}}))$ is the corresponding sine similarity. This adaptive weighting amplifies corrections when the two paths diverge (lower $\alpha_{t_{i}}$, higher $\beta_{t_{i}}$), pulling the result toward the guided direction, and reduces corrections when they agree (higher $\alpha_{t_{i}}$, lower $\beta_{t_{i}}$), preserving the model’s natural prediction.
In score-based (Lipman et al., 2022) terms, $\mathbf{v}_{t_{i}}^{\text{ori}}$ follows the pretrained data manifold, while $\mathbf{v}_{t_{i}}^{\text{traj}}$ may deviate due to injected trajectory signals. DSG applies cosine-weighted interpolation between the two, dynamically combining them to retain trajectory guidance while projecting the velocity back toward the manifold. This improves generation quality by suppressing off-manifold drift. As a result, $\mathbf{v}_{t_{i}}^{\text{corr}}$ drives the sample along the desired motion path while preserving visual fidelity to the model priors. Subsequent experiments confirm the effectiveness of this mechanism.

### v2

@@ 3.1.1 Denoising Diffusion Models and Guidance
Diffusion Solvers. Generative models are largely diffusion (Ho et al., 2020) or flow-based (Lipman et al., 2022). Under the SDE view, diffusion models have a deterministic ODE limit that connects to flow models via reparameterization (Gao et al., 2025) (The detailed derivation is provided in Appendix A). We use the popular DDIM sampler (Song et al., 2020a) as an example to illustrate the sampling process: it recovers the clean sample $\mathbf{x}_{0}$ by reversing the forward noising of a Gaussian prior $\mathbf{x}_{T}$. Given a noise-prediction network $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, the sampler estimates an intermediate signal $\hat{\mathbf{x}}_{0}$ from the current state $\mathbf{x}_{t}$:
@@
Iterating from $t=T$ to $t=0$ yields the final sample $\mathbf{x}_{0}$. Our method intervenes at this stage by modifying $\hat{\mathbf{x}}_{0}$ to enforce trajectory control. Notably, other popular solvers, such as UniPC (Zhao et al., 2023), EDM (Karras et al., 2022), and PNDM (Liu et al., 2022a), also compute $\hat{\mathbf{x}}_{0}$ directly or can recover it via a parameterized transformation, so our framework is broadly compatible. Since our experiments primarily use the flow-based Wan2.1 (Wan et al., 2025) model, we will subsequently detail our algorithm using a flow-based formulation.
Classifier Free Guidance. To improve fidelity to the condition, CFG (Ho & Salimans, 2021) adjusts the network prediction during sampling as:
$\tilde{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},t)=\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})+\omega_{\text{CFG}}\cdot\left[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,\phi)\right],$ (3)
where $\omega_{\text{CFG}}$ is the guidance weight, with $\mathbf{c}$ and $\phi$ denoting the conditional and unconditional inputs, respectively. This interpolates conditional and unconditional scores to steer the sampling trajectory. Our approach extends this principle through a self-referential guidance mechanism that dynamically adjusts the guided prediction using the model’s own unguided output at each step.
@@ 3.1.2 Trajectory Control via Depth-Based Warping
Our framework controls trajectories using a depth-based warping-and-repainting strategy. First, a depth-prediction network estimates camera poses and depth maps $(\mathbf{P}_{q},\mathbf{D}_{q})$ from single image $\mathbf{I}$} or image sequence $\{\mathbf{I}_{q}\}_{q=1}^{N}$ via a function $f:\{\mathbf{I}_{q}\}_{q=1}^{N}\rightarrow\{\mathbf{P}_{q},\mathbf{D}_{q}\}$. Next, a warping operator $\mathcal{W}$ uses these estimates to project a source frame $\mathbf{I}_{src}$ with depth $\mathbf{D}_{src}$ from pose $\mathbf{P}_{src}$ to a target pose $\mathbf{P}_{tar}$. This yields a partial target view $\mathbf{I}^{\prime}_{tar}$ and a validity mask $\mathbf{M}_{tar}$ indicating visible pixels:
@@
The resulting warped frames and masks guide the VDMs along the target poses $\mathbf{P}_{tar}$. This guidance is limited to regions visible in the source views. With these preliminaries, we use this trajectory control to guide video generation in VDMs.
@@ 3.2 Intra-Step Recursive Refinement
To enable precise trajectory injection during VDM’s inference processing, we introduce Intra-Step Recursive Refinement (IRR). As noted in Sec. 3.1.1, the denoising process produces an intermediate variable $\hat{\mathbf{x}}_{0}^{(t)}$, a coarse estimate of the final output and the baseline for later steps,where $t$ denotes the current timestep. IRR modifies $\hat{\mathbf{x}}_{0}^{(t)}$ to impose trajectory constraints, ensuring that generation follows the desired path.
IRR operates within the updates of Eq. (1) and Eq. (2). Given the one-step denoised sample $\hat{\mathbf{x}}_{0}^{(t)}$ from Eq. (1), we fuse it with the trajectory latent $\mathbf{Z}_{\text{traj}}$, obtained by encoding the warped frames of Eq. (4) into latent space. We then add Gaussian noise $\bm{\epsilon}$ to obtain the modified latent $\mathbf{x}_{t}^{\prime}$:
$\mathbf{x}_{t}^{\prime}=(1-w(\sigma))\,\mathbf{F}(\hat{\mathbf{x}}_{0}^{(t)},\mathbf{Z}_{\text{traj}})+w(\sigma)\,\cdot\bm{\epsilon},$ (5)
where $\mathbf{F}(\hat{\mathbf{x}}_{0}^{(t)},\mathbf{Z}_{\text{traj}})=\mathbf{M}\cdot\mathbf{Z}_{\text{traj}}+(1-\mathbf{M})\cdot\hat{\mathbf{x}}_{0}^{(t)}$ copies observable warped content from $\mathbf{Z}_{\text{traj}}$ into the corresponding locations of $\hat{\mathbf{x}}_{0}^{(t)}$ using the binary mask $\mathbf{M}$ from Eq. (4); and $\bm{\epsilon}=\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a randomly sampled Gaussian noise used to re-noise the fused latent $\mathbf{F}(\hat{\mathbf{x}}_{0}^{(t)},\mathbf{Z}_{\text{traj}})$ so that it re-enters the denoising schedule with the injected trajectory. The re-noising is controlled by a scheduler $w(\sigma)$, and the specific strategy can differ for various diffusion and flow models. The re-noised latent $\mathbf{x}_{t}^{\prime}$ is then fed into the network $\bm{\epsilon}_{\theta}$, replacing the original $\mathbf{x}_{t}$ in Eq. (1) and Eq. (2) for the next sampling step. In summary, IRR embeds a micro predictor–corrector at each denoising step. By updating $\hat{\mathbf{x}}_{0}^{(t)}$ with explicit trajectory cues, it continually corrects the sampling path and ensures that synthesis follows the target trajectory precisely.
@@ 3.3 Flow-Gated Latent Fusion
In the IRR process, overwriting all latent channels with trajectory information degrades visual quality because VAE latents encode both motion and appearance. The indiscriminate update in Eq. (5) injects noise into appearance-focused channels. To address this, we propose Flow-Gated Latent Fusion (FLF), a method that identifies and updates latent channels with high motion relevance.
To select motion-related channels, we use an optical-flow-based scoring scheme since flow directly reflects inter-frame motion. First, for each channel $c$ of the latent $\hat{\mathbf{x}}_{0}^{(t)}$ at timestep $t$, we compute the optical flow between consecutive frames to get a predicted flow $\mathcal{F}^{(t,c)}_{\text{pred}}$. The resulting map for each channel has a shape of $2,\tau,H,W]$ (flow vectors, frame pairs, spatial dimensions). Second, we compute a GT reference flow $\mathcal{F}^{(t,c)}_{\text{gt}}$ from the target trajectory latent $\mathbf{Z}_{\text{traj}}$ in the same frame-by-frame manner. Finally, we compare the two flows within the visible regions defined by the mask $\mathbf{M}^{(c)}$.
The comparison relies on three popular optical flow metrics (Teed & Deng, 2020): Masked End-point Error (M-EPE), which measures the Euclidean distance between predicted and GT flow vectors; Masked Angular Error (M-AE), which measures their angular difference; and Outlier Percentage (Fl-all), which calculates the fraction of unreliable pixels. We combine the normalized metrics to calculate a motion similarity score $S^{(t,c)}$ for each channel in each timestep $t$:
$S^{(t,c)}=\sum_{k\in\{\text{E, A, F}\}}\gamma_{k}\!\left(1-{\text{Norm}_{k}}^{(t,c)}\right),$ (6)
where $\text{Norm}_{k}^{(t,c)}\!\in[0,1]$ are the normalized errors from M-EPE, M-AE, and Fl-all, and $\gamma_{k}$ are the weighting factors. The detailed calculation and normalization methods for these metrics are provided in the Appendix C. Higher $S^{(t,c)}$ means better flow alignment and stronger motion evidence.
To adaptively set the motion similarity threshold for each scene, we select motion-relevant channels using a dynamic threshold $\delta^{(t)}=\mu_{S}^{(t)}-\lambda^{(t)}\sigma_{S}^{(t)}$, where $\mu_{S}^{(t)}$ and $\sigma_{S}^{(t)}$ are the mean and standard deviation of all channel scores at step $t$. We schedule $\lambda^{(t)}$ to create a loose-to-tight selection over time. This matches the generative process, where early steps define broad structures and later steps handle fine details. Consequently, we guide more channels initially for structural integrity and fewer channels toward the end to preserve high-fidelity details.
Finally, the latent update selectively fuses the trajectory information $\mathbf{Z}_{\text{traj}}$ into the selected high motion-relevance channels:
@@
$\mathbf{FLF}(\hat{\mathbf{x}}_{0}^{(t)},\mathbf{Z}_{\text{traj}})=\begin{cases}\mathbf{M}^{(c)}\cdot\mathbf{Z}_{\text{traj}}^{(c)}+\left(1-\mathbf{M}^{(c)}\right)\cdot\hat{\mathbf{x}}_{0}^{(t,c)},&\text{if}\;S^{(t,c)}\geq\delta^{(t)}\\ \hat{\mathbf{x}}_{0}^{(t,c)},&\text{otherwise}.\end{cases}$ (7)
This FLF operator replaces the global update in Eq. (5), resulting in a more precise fusion rule: This FLF operator replaces the operator $\mathbf{F}$ in Eq. (5), resulting in a more precise fusion rule.
In summary, FLF provides fine-grained trajectory control while preserving model priors and synthesis quality. Unlike methods that restart the sampling schedule (Xu et al., 2023) or globally update the entire latent (Liu et al., 2024), FLF integrates with our IRR framework to apply selective, per-step guidance, effectively decoupling motion and appearance for precise control.
@@ 3.4 Dual-Path Self-Corrective Guidance
Trajectory latents $\mathbf{Z}_{\text{traj}}$ obtained by warping often contain distortions from depth errors, occlusions, or misalignments, which degrades synthesis quality. To mitigate this, we draw inspiration from CFG (Ho & Salimans, 2021). Conceptually, in the context of a flow model, CFG treats the unconditional prediction as a “bad” direction $\mathbf{v}_{\text{uncon}}$ and the conditional one as a “good” direction $\mathbf{v}_{\text{con}}$ (Karras et al., 2024). It then finds a “better” direction by pushing the “good” one away from the “bad” one. Based on this idea, we propose Dual-Path Self-Corrective Guidance (DSG). At each iteration, IRR produces two velocity fields. The unguided velocity $\mathbf{v}_{t}^{\text{ori}}$ (from the original latent $\mathbf{x}_{t}$) stays on the data manifold with high fidelity but ignores the trajectory, which we consider a “bad” direction for control. The guided velocity $\mathbf{v}_{t}^{\text{traj}}$ (from the corrected latent $\mathbf{x}^{\prime}_{t}$) may be noisy but follows the trajectory, which we consider a “good” direction. DSG uses the difference between them to find a “better” path.
However, we emperically find the difference between our $\mathbf{v}_{t}^{\text{traj}}$ and $\mathbf{v}_{t}^{\text{ori}}$ is far greater than that between the $\mathbf{v}_{\text{con}}$ and $\mathbf{v}_{\text{uncon}}$ in standard CFG. In extensive tests, we observed that the cosine similarity $\alpha_{t}=(\mathbf{v}_{t}^{\text{traj}}\cdot\mathbf{v}_{t}^{\text{ori}})/(\lVert\mathbf{v}_{t}^{\text{traj}}\rVert\cdot\lVert\mathbf{v}_{t}^{\text{ori}}\rVert)$ between our two paths is typically between 0.3–0.6 (an angle of 50°–70°). In contrast, the cosine similarity between $\mathbf{v}_{\text{con}}$ and $\mathbf{v}_{\text{uncon}}$ is nearly 1 (an angle close to 0°), an observation consistent with reports by Wan et al. (2025). Therefore, directly applying the CFG formula fails in our case. To address this large angular difference, we modify the guidance formula to only use the component of the “good” direction that is orthogonal to the “bad” direction. This is achieved by projecting $\mathbf{v}_{t}^{\text{traj}}$ onto $\mathbf{v}_{t}^{\text{ori}}$ (after normalizing them to equal length) and taking the difference, which avoids the adverse effects of their large directional divergence:
@@
$\mathbf{v}_{t}^{\text{corr}}=\mathbf{v}_{t}^{\text{traj}}+\rho\cdot\beta_{t}\bigl(\mathbf{v}_{t}^{\text{traj}}-\alpha_{t}\cdot\mathbf{v}_{t}^{\text{ori}}\bigr),$ (8)
where $\rho$ controls guidance strength, $\alpha_{t}$ is the cosine similarity, and $\beta_{t}=\sqrt{1-\alpha_{t}^{2}}$ is its sine counterpart. The role of $\beta_{t}$ is to adaptively scale the guidance: it amplifies the correction when the paths diverge (low $\alpha_{t}$, high $\beta_{t}$) and reduces it when they agree, preserving the model’s natural prediction.
In essence, DSG robustly balances precise trajectory control with high visual fidelity. Its adaptive cosine-weighting is crucial for handling the large angular difference between the guided and unguided paths. It suppresses artifacts by applying strong corrections when the paths diverge, while preserving the model’s natural predictions when they align. This ensures the final corrected velocity, $\mathbf{v}_{t}^{\text{corr}}$, steers the sample along the target motion path while maintaining the quality of the model’s priors. The effects and details of DSG are presented in Appendix D.